ISRAEL JOURNAL OF MATHEMATICS

TABLE OF CONTENTS, VOLUME 217, 2017

Bardestani, M. and K. Mallahi-Karai On a generalization of the Hadwiger–Nelson problem

313

Björklund, M. Random walks on countable groups  371

Boggarapu, P. and S. Thangavelu On the chaotic behavior of the Dunkl heat semigroup on weighted L^{p} spaces

57

Caicedo, A. E., P. Larson, G. Sargsyan, R. Schindler, J. Steel and M. Zeman Square principles in ℙ_{max} extensions

231

Cuny, C. and M. Weber Ergodic theorems with arithmetical weights

139

Dalla Piazza, F., A. Fiorentino and R. Salvati Manni Plane quartics: the universal matrix of bitangents

111

Fiorentino, A. See F. Dalla Piazza Funano, K. Estimates of eigenvalues of the laplacian by a reduced

number of subsets

413

Hempel, N. and A. Onshuus Groups in NTP_{2}  355

Hughes, K. Maximal functions and ergodic averages related to Waring’s problem

17

Kolman, O. and B. Wald M-slenderness  303

Kuznetsvova, Y. and J. Roydor Homomorphisms with small bound between Fourier algebras

283

Lacey, M. T. An elementary proof of the A_{2} bound  181
Lambert, V., T. H. Lê and A. Plagne Additive bases in groups  383

Larson, P. See A. E. Caicedo Lê, T. H. See V. Lambert Mallahi-Karai, K. See M. Bardestani Onshuus, A. See N. Hempel

Pakovich, F. On decompositions of trigonometric polynomials  337

Pérez, C. and I. P. Rivera-Ríos Borderline weighted estimates for commutators of singular integrals

435

Plagne, A. See V. Lambert Rivera-Ríos, I. P. See C. Pérez Roydor, J. See Y. Kuznetsvova Salvati Manni, R. See F. Dalla Piazza Sargsyan, G. See A. E. Caicedo Schindler, R. See A. E. Caicedo Schlage-Puchta, J.-C. and G. Weitze-Schmithüsen Finite

translation surfaces with maximal number of translations

1

Schröder, J. P. Minimal rays on closed surfaces  197

Shapira, A. and R. Yuster A tournament approach to pattern avoiding matrices

477

Siciliano, S. and H. Usefi Lie structure of smash products  93

Steel, J. See A. E. Caicedo Thangavelu, S. See P. Boggarapu Usefi, H. See S. Siciliano Wald, B. See O. Kolman Weber, M. See C. Cuny Weitze-Schmithüsen, G. See J.-C. Schlage-Puchta Wise, D. T. and D. J. Woodhouse A cubical flat torus theorem

and the bounded packing property

263

Woodhouse, D. J. See D. T. Wise Yuster, R. See A. Shapira Zeman, M. See A. E. Caicedo